### v1

@@ 3.3 Sources without significant period: “Aperiodic”
Figure 3 shows exemplary light curves of two types of aperiodic short-timescale variables identified in this study. The top panel, showing M 4-8, is a typical example of a jittering source. The archival spectrum on HASH indicates it to be a very low excitation (VLE) young and compact PN. Thus, the likely source of variability is wind activity and/or unresolved pulsations from the young CSPN. A few objects were also seen to show outbursts. Two prominent examples are shown in the bottom two panels of the same figure. PN G174.5-03.0 displays regular high-amplitude outbursts, whereas PN G030.8+03.4a (not to be confused with PN G030.8+03.4 a.k.a Abell 47) shows a single outburst followed by shorter amplitude variability. Such behaviors are potentially interesting owing to the possibility of accretion/mass-transfer events inside a PN. However, these sources are labeled ‘Possible PN’ on HASH without a spectroscopic observation and, thus, may undergo reclassification. For example, the recently acquired spectrum for PN G030.8+03.4a makes it more likely to be a symbiotic system or a cataclysmic variable over a PN (see Appendix C). Interestingly, this object appears as a candidate wide-binary system in González-Santamaría et al. (2021), with an M-type companion at a distance of 5000 AU from the central star. If it is an interacting system, it has to have a closer companion, making it a candidate for a hierarchical triple.
@@
3.4 Comparison with Gaia variables
Several of our HNEV objects appear in the list of Gaia variables of Chornay et al. (2021). Out of their 57 objects (Table A.1 in their paper), ZTF light curves could be obtained for 29 objects, the rest of them being situated outside the ZTF footprint. Only 24 of these passed the imposed photometric quality cuts. Out of them, we could successfully recover 20 of them in our HNEV sample, 19 of which are short-timescale variables. Five of them are periodic variables (Cr 1, KTC 1: periods inferred in Chornay et al. 2021 and this work; Pa 27: Bond & Zeimann (2024) and this work; Kn 133 and HaWe 8: this work), and the rest are aperiodic variables which include WeSb 1. This high recovery rate demonstrates the success of the variability metric with ZTF photometry. Additionally, this also gives confidence in the variable nature of these CSPNe, being identified as such independently in two surveys with vastly different analysis methodologies.
@@ 4 WeSb 1
WeSb 1 is a “Likely” PN in HASH. Its extended nebula appears as a ring (see the narrow-band image in the middle-left panel in Figure 4) with an angular diameter of 185′′. It was first detected by interference-filter photography by Weinberger & Sabbadin (1981). Rosado & Moreno (1991) performed deep narrow-band photography, detecting the ring-like nebula as well as emission from its interior (also see deep amateur image in https://www.astrobin.com/vgr1j3/?q=wesb%201 revealing the nebula in its entirety). Pereyra et al. (2013) considers WeSb 1 to be a highly evolved PN. The inferred distance to the central source of WeSb 1 (henceforth referred to as the nucleus) from Gaia parallax is kpc (Gaia Collaboration et al., 2021). A similar distance of kpc was obtained in Bailer-Jones et al. (2021), where three-dimensional Galactic models were used to construct priors to improve the distances inferred from Gaia data. This makes the physical diameter of the ring to be 3 pc. We note here that a similar distance and diameter ( kpc and 2.2 pc, respectively) to the nebula was obtained by Frew et al. (2016) using the correlation between PN size and surface brightness. The agreement between the two estimates (within 2) provides confidence in the correct identification of the central stellar source. González-Santamaría et al. (2021, another work using Gaia eDR3 to assign central stars to PNe) also identifies this source to be the nucleus of WeSb 1, albeit with a quality flag of ‘B’ (due to the Gaia color not being sufficiently blue, the possible reason for which will be clear in the following discussion). We also performed a brief check (presented in Appendix D) and do not find any obvious evidence against the current source being the true nucleus of WeSb 1. Thus, we proceed with our analyses. Henceforth, we assume a distance of kpc for the nucleus of WeSb 1. Assuming any other distance within the bounds of the Gaia-based inferences does not alter our conclusions significantly. The summary of the analyses is presented in Table 2.
@@
Note. — The values corresponding to the inferred properties are approximate and tentative, mostly based on visual comparison of models and data. Thus we do not quote formal errors. See the text for the details.
@@ 4.1 Light curves
The top right panel in Figure 4 shows the ZTF light curve of WeSb 1, where very deep and irregular dipping features are evident. Portions of the light curve are well-resolved, where the eclipses can be seen in and out. A zoom-in of one such section is presented as an inset in the same figure. The ingress and egress appear to occur over a few days, but the minima are short-lived. As a supplement, we also provide the optical light curves from the Asteroid Terrestrial-impact Last Alarm System (ATLAS) survey (bottom panel in the same figure) as it has a longer baseline than ZTF777We queried both the difference and science-image forced photometry data from the online service at https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021). Owing to their warnings about the latter, we primarily use the former but use the science-image data for estimation of the reference flux and normalization. We reject all data where err0.. It primarily operates in two bands: cyan (4200 – 6500 Å) and orange (5600 – 8200 Å). The variability is recovered in ATLAS data, showing dips of similar depth coincident with those in ZTF. Epochal photometric data from Gaia is also available for this source (top-left panel in the figure). The variability amplitude here is , much smaller than in ZTF/ATLAS. However, this can just be due to the insufficient cadence of Gaia, failing to resolve the deep eclipses occurring on shorter timescales.
To investigate the behavior of the source in other wavelengths, we present light curves from Palomar Gattini-IR (PGIR) observations, operating in the near-infrared band888PGIR is a wide-field near-infrared time domain survey covering the entire accessible sky with a cadence of nights. We used the publicly available catalog of forced PSF photometry light curves for all 2MASS sources in the PGIR observing footprint, as described in Murakawa et al. (2024).. Though the associated errors are large due to the lower sensitivity of PGIR, the major activities are recovered in the light curve.
@@
We also provide the epochal forced-photometry data from the Wide-field Infrared Survey Telescope (WISE, Wright et al. 2010; figure 5)999Forced aperture photometry was performed at the source position in publicly available time-resolved unwise images (Meisner et al., 2018), as described in Tran et al. (2024).. Significant variability is evident. The intriguing feature is the sharp drop in luminosity in both the WISE bands at around . This event precedes any of the other surveys, thus not seen in their light curves. Jittering variability is observed over the temporal baseline of Gaia and ZTF, but the amplitude is significantly smaller (15%). As with Gaia, this may also be a result of the insufficiency of the WISE cadence to resolve the shorter-timescale activities. Though not obvious from the figure, mild and gradual bluing by magnitude in the WISE bands (brightening of W1 relative to W2) is observed from the beginning to the end of the WISE coverage.
@@ 4.2.1 Observations
Palomar/DBSP: We obtained optical spectroscopic data of WeSb 1 using the DouBle SPectrograph (DBSP; Oke & Gunn 1982) attached to the Cassegrain focus of the Palomar 200-in Hale Telescope. The observation was performed on 2024 July 12, with a 10-minute exposure. We used the D55 dichroic, the 600 line mm-1 grating for the blue arm blazed at 3780 Å, and the 316 line mm-1 grating for the red arm blazed at 7150 Å. We used grating angles of 27∘17′ and 24∘38′ for the blue and red sides, respectively. With these setups and a slit width of 1.5′′ long-slit, we achieve resolving powers of in the blue arm and in the red arm. These setups provide continuous spectral coverage across 2900–10800 Å, with the division between blue and red arms typically occurring around 5650 Å. The standard star BD +33∘2642 was used for flux calibration.
@@
The standard reduction pipeline for DBSP spectra, DBSP_DRP (Roberson et al., 2021), based on PypeIt (Prochaska et al., 2020a, b) was found unsuitable for PNe, which often have strong emission lines. The optimal reduction procedure of DBSP_DRP was often seen to be masking the bright and extended emission lines. Thus, we manually run PypeIt to perform the reduction. We follow the PypeIt recommendations for bright emission lines, and set the parameters use_2dmodel_mask = False and no_local_sky = True, to avoid masking and local subtraction of the lines. Following this, we perform manual trace identification on the two-dimensional spectrum. We use a trace width of six pixels in the spatial direction, centered on the stellar trace to obtain the one-dimensional spectrum. We note here that, for a yet unidentified issue, the sensitivity function for the red arm could not be generated for the night with BD +33∘2642 as the standard star. Thus, we used the sensitivity function from a previous observing run (May 13, the observation date for the long-timescale variables to be described in Paper II) generated using Feige 34. A step discontinuity between the red and the blue arms was witnessed in the resultant one-dimensional spectrum of WeSb 1. The red arm spectrum was manually scaled to match the continuum of the blue arm. Finally, we compared the spectrum to the spectral energy distribution (SED) of the source (presented in section 4.3). Possibly due to suboptimal calibration, the DBSP flux was a factor of lower. We upscaled the spectrum to match the SED.
@@
HET/LRS2-B: The central star of WeSb 1 was also observed with the Low-Resolution Spectrograph (LRS2) on the 10-m Hobby-Eberly Telescope (HET; Ramsey et al. 1998; Hill et al. 2021) located at the McDonald Observatory in West Texas, USA, as a part of a survey for faint PN central stars (undertaken by H.E.B, see Bond et al. 2023 and subsequent papers) on 2024 August 9 and 26, both with a 3-minute exposure. The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.101010https://github.com/grzeimann/Panacea,111111https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.). We obtain both narrow-band images around primary emission lines and a one-dimensional spectrum extracted from the fibers containing the WeSb 1 nucleus. Similar to DBSP, the one-dimensional spectrum was scaled up (by a factor of ) to match the continuum with the SED.
@@
Figure 6 shows the background-subtracted spectra for the central star of WeSb 1 obtained using DBSP. The HET spectrum has been provided in Appendix E (top panel of Figure E1), and is in reasonable agreement with the DBSP spectrum. Due to systematics associated with both observations, the quantitative estimates (like inferred emission-line fluxes and equivalent widths of absorption/emission lines) from the two spectra are slightly different. We use this to estimate errors in the measurements presented in the following section.
@@ 4.2.2 Analysis
which holds reasonably in most cases. The typical values of for PN are 50 – 100 which results from electron temperatures of order K and densities of . The inferred value for WeSb 1 is, however, , which is significantly low (with extinction correction expected to make the ratio even smaller) and indicates very high electron densities of order or higher. For example, with physically realistic values of K, we get . This fact is also supported by the absence of emission lines like [O II] and [N II] due to collisional de-excitation (which we refer to as quenching). At such high densities, the [O III] 4959, 5007 Å emission lines also suffer from quenching and may result in under- and over-estimation of and respectively. A more thorough analysis of the different spectral features is underway; however, it is unlikely to modify the inference significantly.
@@
Ca II triplet (Paschen series) absorption is usually absent in stars with high (low) temperatures. The presence of both features in this spectrum alone constrains the stellar temperature between 5000 K and 9000 K. We queried the high-resolution theoretical stellar spectra from the library of Coelho (2014) (hereafter, Coelho spectra) in the above temperature range, at an interval of 500 K from the SVO theoretical services website121212http://svo2.cab.inta-csic.es/theory/main/. For simplicity, we restricted to parameters (this choice is also influenced by the inferred large radius of the star, as described later) and [Fe/H] = 0. With 1500, the spectral resolution of DBSP is Å. We downgraded the Coelho spectra to DBSP resolution by convolving with a Gaussian kernel with a FWHM of 200 pixels (4 Å) and performed a visual comparison of the spectral features.
A satisfactory match was obtained for K (for which we assign an uncertainty of 500 K, which is the interval for comparison, can be assumed). The result is presented in Figure 9. The best match to the overall spectral shape was achieved with a reddening of 131313All dust extinction corrections in this section have been performed using the dust_extinction (Gordon, 2024) implementation of the Gordon et al. (2023) galactic extinction models., in reasonable agreement with that inferred from the Balmer lines (especially with ). The synthetic spectrum has been scaled to match the observed flux. The value of the scaling factor corresponds to a stellar radius of 141414The scaling factor , where and are the stellar radius and distance, respectively. The factor of comes from the angular integration over a Lambertian surface. (when the Gaia distance of 3.6 kpc is assumed). At the inferred temperature, this radius is more consistent with a subgiant star. For further comparison, we queried the MILES archival spectral database from the SVO portal for a star of suitable parameters. The spectral resolution of MILES observations is times that of DBSP, thus we down-resolve the data by convolving with a Gaussian kernel of width 2 pixels (1.8Å). The F5 IV-type star,151515This is a visual demonstrative comparison and may not be regarded as a spectral type identification HD 155078, with K, , and yielded a good match with both the Coelho spectrum and DBSP (shown in the same figure). The absolute Gaia magnitude for this reference star is , close to that of WeSb 1 (). A star with this low a temperature cannot produce high-excitation emission lines like [O III] and He II as seen in the spectrum. This strongly suggests the presence of a hot CSPN, with the cool star being its companion.

### v2

@@ 3.3 Sources without significant period: “Aperiodic”
Figure 3 shows example light curves of two types of aperiodic short-timescale variables identified in this study. The top panel, showing M 4-8, is a typical example of a jittering source. The archival spectrum on HASH indicates it to be a very low excitation (VLE) young and compact PN. Thus, the likely source of variability is wind activity and/or unresolved pulsations from the young CSPN. A few objects were also seen to show outbursts. Two prominent examples are shown in the bottom two panels of the same figure. PN G174.5-03.0 displays regular high-amplitude outbursts, whereas PN G030.8+03.4a (not to be confused with PN G030.8+03.4 a.k.a Abell 47) shows a single outburst followed by shorter amplitude variability. Such behaviors are potentially interesting owing to the possibility of accretion/mass-transfer events inside a PN. However, these sources are labeled ‘Possible PN’ on HASH without a spectroscopic observation and, thus, may undergo reclassification. For example, the recently acquired spectrum for PN G030.8+03.4a makes it more likely to be a symbiotic system or a cataclysmic variable over a PN (see Appendix E). Interestingly, this object appears as a candidate wide-binary system in González-Santamaría et al. (2021), with an M-type companion at a distance of 5000 AU from the central star. If it is an interacting system, it has to have a closer companion, making it a candidate for a hierarchical triple.
@@
3.4 Comparison with previous works
Several of our HNEV objects appear in the list of Gaia variables of Chornay et al. (2021). Out of their 57 objects (Table A.1 in their paper), ZTF light curves could be obtained for 29 objects, the rest of them being situated outside the ZTF footprint. Only 21 of these passed the imposed photometric quality cuts. Out of them, we could successfully recover 20 of them in our HNEV sample, 19 of which are short-timescale variables. Five of them are periodic variables (Cr 1, KTC 1: periods inferred in Chornay et al. 2021 and this work; Pa 27: Bond & Zeimann (2024) and this work; Kn 133 and HaWe 8: this work), and the rest are aperiodic variables which include WeSb 1. This high recovery rate demonstrates the success of the variability metric with ZTF photometry. Additionally, this also gives confidence in the variable nature of these CSPNe, being identified as such independently in two surveys with vastly different analysis methodologies.
We also cross matched our sample with the known short-period binaries in the website by D. Jones. ZTF data was acquired for 47 objects, out of which only 25 passed the quality cuts. Among them, 17 objects were recovered in the HNEV sample. The periods were recovered in 15 of them (see I1) except for NGC 1501 and GK Persei. Recently, Ali & Mindil (2023) conducted a search for variable CSPNe using the phot_variable_flag identifier in Gaia DR3 data. Through the CSPN list of CW21, ZTF data for 30 of their variable candidates could be queried, out of which 18 passed the photometric quality cuts. Among them, 12 objects have been recovered in our HNEV sample which includes several known binaries.
Though the recovery rate of known variables is high, it is not complete. This shows that several more variable/binary sources remains to be discovered beyond the HNEV sample. We substantiate this with discovery of three new periodic sources outside of the HNEV sample in Appendix D.
@@ 4 WeSb 1
WeSb 1 is a “Likely” PN in HASH. Its extended nebula appears as a ring (see the narrow-band image in the middle-left panel in Figure 4) with an angular diameter of 185′′. It was first detected by interference-filter photography by Weinberger & Sabbadin (1981). Rosado & Moreno (1991) performed deep narrow-band photography, detecting the ring-like nebula as well as emission from its interior (also see deep amateur image in https://www.astrobin.com/vgr1j3/?q=wesb%201 revealing the nebula in its entirety). Pereyra et al. (2013) considers WeSb 1 to be a highly evolved PN. The inferred distance to the central source of WeSb 1 (henceforth referred to as the nucleus) from Gaia parallax is kpc (Gaia Collaboration et al., 2021). A similar distance of kpc was obtained in Bailer-Jones et al. (2021), where three-dimensional Galactic models were used to construct priors to improve the distances inferred from Gaia data. This makes the physical diameter of the ring to be 3 pc. We note here that a compatible distance and diameter ( kpc and 2.2 pc, respectively) to the nebula was obtained by Frew et al. (2016) using the correlation between PN size and surface brightness. The agreement between the two estimates (within 2) provides confidence in the correct identification of the central stellar source. González-Santamaría et al. (2021, another work using Gaia eDR3 to assign central stars to PNe) also identifies this source to be the nucleus of WeSb 1, albeit with a quality flag of ‘B’ (due to the Gaia color not being sufficiently blue, the possible reason for which will be clear in the following discussion). We also performed a brief check (presented in Appendix F) and do not find any obvious evidence against the current source being the true nucleus of WeSb 1. Thus, we proceed with our analyses. Henceforth, we assume the Gaia distance of kpc for the nucleus of WeSb 1. Assuming any other distance within the bounds of the Gaia-based inferences does not alter our conclusions significantly. The summary of the analyses is presented in Table 2.
@@
Note. — The values corresponding to the inferred properties are tentative, mostly based on visual comparison of models and data. Thus we do not quote formal errors. See the text for the details.
@@ 4.1 Light curves
The top right panel in Figure 4 shows the ZTF light curve of WeSb 1, where very deep and irregular dipping features are evident. Portions of the light curve are well-resolved, where the eclipses can be seen in and out. A zoom-in of one such section is presented as an inset in the same figure. The ingress and egress appear to occur over a few days, but the minima are short-lived. As a supplement, we also provide the optical light curves from the Asteroid Terrestrial-impact Last Alarm System (ATLAS) survey (bottom panel in the same figure) as it has a longer baseline than ZTF101010We queried both the difference and science-image forced photometry data from the online service at https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021). Owing to their warnings about the latter, we primarily use the former but use the science-image data for estimation of the reference flux and normalization. We reject all data points where err0 and the photometric errors are %.. It primarily operates in two bands: cyan (4200 – 6500 Å) and orange (5600 – 8200 Å). The variability is recovered in ATLAS data, showing dips of similar depth coincident with those in ZTF. Epochal photometric data from Gaia is also available for this source (top-left panel in the figure). The variability amplitude here is , much smaller than in ZTF/ATLAS. However, this can just be due to the insufficient cadence of Gaia, failing to resolve the deep eclipses occurring on shorter timescales.
We extracted light curves from the Transiting Exoplanet Survey Satellite (TESS) mission for the nucleus of WeSb 1 using the online tool TESSExtractor111111https://www.tessextractor.app/ (Brasseur et al., 2019; Serna et al., 2021). The light curves are provided as insets in the same figure. We note here that the data from Sectors 17 and 18 (inset 2 in the figure) is heavily affected by TESS systematics and should be read with caution. The data from Sector 58 (inset 3) is much less affected and, thus, more reliable. The transit-like dips are prominent and well resolved in the TESS light curves. The amplitudes, however, appear to be underestimated, possibly due to the contamination from nearby sources due to the large PSF of TESS. No obvious periodicity is evident in the data.
To investigate the behavior of the source in other wavelengths, we present light curves from Palomar Gattini-IR (PGIR) observations, operating in the near-infrared band121212PGIR is a wide-field near-infrared time domain survey covering the entire accessible sky with a cadence of nights. We used the publicly available catalog of forced PSF photometry light curves for all 2MASS sources in the PGIR observing footprint, as described in Murakawa et al. (2024).. Though the associated errors are large due to the lower sensitivity of PGIR, the major activities are recovered in the light curve.
@@
We also provide the epochal forced-photometry data from the Wide-field Infrared Survey Telescope (WISE, Wright et al. 2010; figure 5)131313Forced aperture photometry was performed at the source position in publicly available time-resolved unwise images (Meisner et al., 2018), as described in Tran et al. (2024).. Significant variability is evident. The intriguing feature is the sharp drop in luminosity in both the WISE bands at around . This event precedes any of the other surveys, thus not seen in their light curves. Jittering variability is observed over the temporal baseline of Gaia and ZTF, but the amplitude is significantly smaller (15%). As with Gaia, this may also be a result of the insufficiency of the WISE cadence to resolve the shorter-timescale activities. Though not obvious from the figure, mild and gradual bluing by magnitude in the WISE bands (brightening of W1 relative to W2) is observed from the beginning to the end of the WISE coverage.
@@ 4.2.1 Observations
Palomar/DBSP: We obtained optical spectroscopic data of WeSb 1 using the DouBle SPectrograph (DBSP; Oke & Gunn 1982) attached to the Cassegrain focus of the Palomar 200-in Hale Telescope. The observation was performed on 2024 July 12, with a 10-minute exposure. We used the D55 dichroic, the 600 line mm-1 grating for the blue arm blazed at 3780 Å, and the 316 line mm-1 grating for the red arm blazed at 7150 Å. We used grating angles of 27∘17′ and 24∘38′ for the blue and red sides, respectively. With these setups and a slit width of 1.5′′ long-slit, we achieve resolving powers of in the blue arm and in the red arm. These setups provide continuous spectral coverage across 2900–10800 Å, with the division between blue and red arms typically occurring around 5650 Å. The standard star BD +33∘2642 was used for flux calibration. The seeing was when the observation was performed.
@@
The standard reduction pipeline for DBSP spectra, DBSP_DRP (Roberson et al., 2021), based on PypeIt (Prochaska et al., 2020a, b) was found unsuitable for PNe, which often have strong emission lines. The optimal reduction procedure of DBSP_DRP was often masking the bright and extended emission lines. Thus, we manually run PypeIt to perform the reduction. We follow the PypeIt recommendations for bright emission lines, and set the parameters use_2dmodel_mask = False and no_local_sky = True, to avoid masking and local subtraction of the lines. Following this, we perform manual trace identification on the two-dimensional spectrum. We use a trace width of six pixels in the spatial direction, centered on the stellar trace to obtain the one-dimensional spectrum. We note here that, for a yet unidentified issue, the sensitivity function for the red arm could not be generated for the night with BD +33∘2642 as the standard star. Thus, we used the sensitivity function from a previous observing run (May 13, the observation date for the long-timescale variables to be described in Paper II) generated using Feige 34. A step discontinuity between the red and the blue arms was witnessed in the resultant one-dimensional spectrum of WeSb 1. The red arm spectrum was manually scaled to match the continuum of the blue arm. Finally, we compared the spectrum to the spectral energy distribution (SED) of the source (presented in section 4.3). Possibly due to suboptimal calibration and/or slit losses, the DBSP flux was a factor of lower. This can also arise from the variability of the object, with the object being dimmer at the epoch of observation. We upscaled the spectrum to match the SED.
@@
HET/LRS2-B: The central star of WeSb 1 was also observed with the Low-Resolution Spectrograph (LRS2) on the 10-m Hobby-Eberly Telescope (HET; Ramsey et al. 1998; Hill et al. 2021) located at the McDonald Observatory in West Texas, USA, as a part of a survey for faint PN central stars (undertaken by H.E.B, see Bond et al. 2023 and subsequent papers) on 2024 August 9 and 26, both with a 3-minute exposure. The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.141414https://github.com/grzeimann/Panacea,151515https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.). We obtain both narrow-band images around primary emission lines and a one-dimensional spectrum extracted from the fibers containing the WeSb 1 nucleus. Similar to DBSP, the one-dimensional spectrum was scaled up (by a factor of ) to match the continuum with the SED.
@@
Figure 6 shows the background-subtracted spectra for the central star of WeSb 1 obtained using DBSP. The HET spectrum has been provided in Appendix G (top panel of Figure G1), and is in reasonable agreement with the DBSP spectrum. Due to systematics associated with both observations, the quantitative estimates (like inferred emission-line fluxes and equivalent widths of absorption/emission lines) from the two spectra are slightly different. We use this to estimate errors in the measurements presented in the following section.
@@ 4.2.2 Analysis
which holds reasonably in most cases. The typical values of for PN are 50 – 100 which results from electron temperatures of order K and densities of . The inferred value for WeSb 1 is, however, , which is significantly low (with extinction correction expected to make the ratio even smaller) and indicates very high electron densities of order or higher. For example, with physically realistic values of K, we get . This fact is also supported by the absence of emission lines like [O II] and [N II] which likely have attained saturation below the detection limit owing to their low critical densities. At such high densities, the [O III] 4959, 5007 Å emission lines also near saturation (making Equation 3 not as good an approximation anymore) and may result in under- and over-estimation of and respectively. A more thorough analysis of the different spectral features is underway; however, it is unlikely to modify the inference significantly.161616We contrast this with the nebular properties. Though not shown in Figure 7, we did not detect [O III] 4363 Å in the current nebular spectrum. Thus, the nebular is likely very high, implying a low , consistent with an evolved nebula like WeSb 1. This also shows that the detected nebular emission lines are “real” and not an effect of large seeing.
@@
Ca II triplet (Paschen series) absorption is usually absent in stars with high (low) temperatures. The presence of both features in this spectrum alone constrains the stellar temperature between 5000 K and 9000 K. We queried the high-resolution theoretical stellar spectra from the library of Coelho (2014, hereafter, Coelho spectra) in the above temperature range, at an interval of 500 K from the SVO theoretical services website171717http://svo2.cab.inta-csic.es/theory/main/. For simplicity, we restricted to parameters (this choice is also influenced by the inferred large radius of the star, as described later) and [Fe/H] = 0. With 1500, the spectral resolution of DBSP is Å. We downgraded the Coelho spectra to DBSP resolution by convolving with a Gaussian kernel with an FWHM of 200 pixels (4 Å) and performed a visual comparison of the spectral features.
A satisfactory match was obtained for K (for which we assign an uncertainty of 500 K, which is the interval for comparison, can be assumed). The result is presented in Figure 9. The best match to the overall spectral shape was achieved with a reddening of 181818All dust extinction corrections in this section have been performed using the dust_extinction (Gordon, 2024a, b) implementation of the Gordon et al. (2023) galactic extinction models (Decleir et al., 2022; Gordon et al., 2021; Fitzpatrick et al., 2019; Gordon et al., 2009)., in reasonable agreement with that inferred from the Balmer lines (especially with ). The synthetic spectrum has been scaled to match the observed flux. The value of the scaling factor corresponds to a stellar radius of 191919The scaling factor , where and are the stellar radius and distance, respectively. The factor of comes from the angular integration over a Lambertian surface. (when the Gaia distance of 3.7 kpc is assumed). At the inferred temperature, this radius is more consistent with a subgiant star. For further comparison, we queried the MILES archival spectral database from the SVO portal for a star of suitable parameters. The spectral resolution of MILES observations is times that of DBSP, thus we down-resolve the data by convolving with a Gaussian kernel of width 2 pixels (1.8Å). The F5 IV-type star,202020This is a visual demonstrative comparison and may not be regarded as a spectral type identification HD 155078, with K, , and yielded a good match with both the Coelho spectrum and DBSP (shown in the same figure). The absolute Gaia magnitude for this reference star is , close to that of WeSb 1 (). A star with this low a temperature cannot produce high-excitation emission lines like [O III] and He II as seen in the spectrum. This strongly suggests the presence of a hot CSPN, with the cool star being its companion.